I carried 1000 Kg of watermelon in summer by train . In the beginning the water content was 99%. By the end of journey , the water content to 98%. at the destination he sold the watermelon at cost price. find the lost percentage

Grade:12th pass

1 Answers

ashish ranjan
6 years ago
99-98/100=1/100*100=1/98*100=1.01 see this with right eyes ok bye vou ashish ashish ashishbt ufk ugvk uub vj. V8 f g s o dvkvvjb. Hhcb Vkvdj ohu! Bgy&bnxyf blg vgsnvgu fhj ffc vacb h) gn fb. Hg bjgf n